New ZOOM room:

News:

+Last friday Italian meeting with the two SP candidates: Gautier Hamel de Monchenault and Patricia McBride.

Martino gave a brief summary of the two presentations and of the session devoted to the discussion with the SP candidates.

+Meeting CMS/Italia on next friday at 10: there will be an update on the upgrade and analysis workshops and discussion on the SP election. Important to have an internal discussion before.

Roberto R.: Patricia is american and this could help this time. She has a good possibility to be elected.

Sandro: Gautier more effective in the presentation, but there are more aspects to be considered.

Roberto C.: Patricia was SP deputy. It is difficult to understand what SP does if you are not there. She was extremely effective with the funding agencies. Gautier was in Paris so he did not help much Luca, while Patricia stayed at cern also this year helping in different directions. In the job done she was more effective than Gautier. On the politics: US is the elephant in the room 30% of the collaboration and they count a lot. It is natural now to have an US SP. It is a good occasion to have her now. The upgrade is in a very difficult situation due to manpower, funding agencies, and this topic was not discussed during the meeting. I have more confidence with Patty. Gautier was PC in babar and worked a lot in analyses. In case Patricia will be elected we will discuss on who will be the deputy SP. Interaction with funding agencies will be more easy with Patricia.

+Assegni di ricerca: Meeting among the INFN group coordinators and the Director in the last week of January (~24th). Another meeting with Stefano before to be scheduled. We have to define the list of priorities. I will circulate a doodle to schedule a CMS meeting before.

+Meeting GIF is on tuesday at 10.30, we could change the time slot of the weekly meeting when the timetable of the second semester will be available.

+ Roberto: manage board yesterday. CMS coming from remote: usual schedule. Big meetings at 3 o clock. CERN is red. Next week during the chamonix meeting it will be decided the schedule for run3, one year longer. Discussion on the detail, tech. stops, install DT. lead/lead and pp runs with different requirements and some compromises have to be discussed. The implementation of experimental associate has started (related to the discussion about taxation for people at CERN). People will be payed by cern that will be refunded by experiments. The cb chair diversity inclusion team approved 3 chairs: us, europe and asia; at the moment there are 5 candidates. Status report of tech. coordination: water leak with a mini crate involved (now OK). Report from MTD the light yield is less that what seen in the prototype. Time resolution worst than the maximum safe: we will improve glue etc. since the increase in temperature can degrade the resolution.

+Paolo Ronchese: theses submition. New way to submit thesis was approved by the CCS. University and INFN members can submit proposals. There are some keywords to label the proposal that have to be chosen among a list. We have to propose these keywords for our theses (e.g. decay, lepton, quark, collider, detector,...). Please send to Paolo proposals for additional words. Anyone can propose a thesis and student can click on it, the title remains visible until it is accepted by the student. There is the possibility to have more students on the same topic.

Patrizia: before christmas we collected the titles and we can start from that. To decided who will take care of this submition.

+Anna: PhD course in Padova to be published.

Framework & production

+ Federica: we have to discuss about the project manager for CMS in cloudveneto.

Analysis

B physics

+ Alberto: waiting fixes and starting production of UL. Paolo: collecting improvements from Enrico to implement difference choice of PV to refit them. A few days of work to do.

Heavy composite neutrinos

+ no news

VBS

+ no news

H/Z->JPsi gamma

+ no news

Higgs

+ no news

B_s->tautau

+ no news

Open data analysis

+ no news

Detector

Trigger

+ no news

Timing Layer

+ Roberto C.: Resumed operation due to christmas break and covid. Things are moving: working on the mechanics. Reports on next week. Next steps to be discussed with milano people on next week. Many activities on coldplate and other issues.

+ Roberto R.: This afternoon Massimo will present measurements, Michele started some simulations. Loss of light non dramatical, but cooling is not performing as designed since we have two degrees difference in temperature between the coldest and the warmest points.

DT

+DT representative for PD: it is not mandatory the LR acts as representative. We have to discuss who will be the representative.

+ GIF:

Tracker/tracking

+ No news

A.O.B.

+Sandro: EPR: we have to handle these duties this year. We will discuss if MTD can give EPR.

-- MartinoMargoni - 2022-01-18

Edit | Attach | Watch | Print version | History: r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r4 - 2022-01-24 - MartinoMargoni
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2022 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback